OpenStudy (anonymous):

Tammy is planting a garden that is 60ft.square feet. How many square inches is the garden?

OpenStudy (anonymous):

How many square inches are in a square foot?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

there are twelve inches along each edge, so how many little squares will you have in the big square?

OpenStudy (anonymous):

144

OpenStudy (anonymous):

Correct. So if there are 144 square inches in each square foot, and you have 60 square feet, how many square inches will you have?

OpenStudy (anonymous):

144x60=8640?

OpenStudy (anonymous):

Also correct

OpenStudy (anonymous):

Thanks that helped a lot!=)

OpenStudy (anonymous):

You can check your answer using unit analyses: \[60ft^2 = 60\times ft \times ft \times \frac{12in}{ft} \times \frac{12in}{ft} = 8640 \times in \times in = 8640 in^2\]

OpenStudy (anonymous):

can you help with my other one?

OpenStudy (anonymous):

I still dont get it